Fermentation is the process of making wine by converting glucose (C6H12O6) into Ethanol(C2H5OH) and carbon dioxide (CO2). How many grams of carbon dioxide would be formed from 26 grams of glucose?
Answer #1

Equation of the reaction : C6H12O6\rightarrow 2 C2H5OH + 2 CO2

Moles of glucose = mass / molar mass = 26 / 180 = 0.1444 moles

From the equation, moles of CO2 = 2 * moles of glucose = 2 * 0.1444 = 0.2888 moles

Mass of CO2 = moles * molar mass = 0.2888 * 44 = 12.71 g